# Probing dielectric breakdown in Mott insulators through current oscillations

### v1

@@ 2.2 Linearly increasing magnetic flux
The non-equilibrium current is shown in Figure 3 with solid lines, while the persistent current is plotted in dotted lines. For $U=0$ the time-evolved wave function is at all times fully populating the same instantaneous eigenstate. As the eigenstate does not correspond, in general, to the ground state, the current does not show the Aharonov period and shows no longer a sawtooth shape. In this case, it oscillates with the Bloch period and has higher amplitude than in equilibrium, as shown in Figure 3a).
Figure 3: Charge current of $L=10$ half-filled rings in the a) non-interacting $(U=0)$ and b) interacting $(U=2)$ regimes. Different curves correspond to different fields. The persistent current is displayed with dotted lines. A similar plot of the interacting regime can be found in [46].
For $U>0$, the situation is different. The existence of the charge gap allows the wave function to remain in the insulating state, as long as the evolution is adiabatic. That is why for sufficiently small fields the current is the same as in equilibrium, as shown in Figure 3b). For stronger electric fields, tunnelling to the excited spectrum becomes relevant, and the persistent current pattern is lost around the first avoided crossing, at $E\cdot t=0.5$. This is a signature of the dielectric breakdown of the system. It is relevant to note that the system does not necessarily excite to the first excited state, since there are excitations that do not couple to the electric field. For even larger fields, a Bloch oscillation pattern is observed, since the wave function can cross multiple avoided crossings and reach the top of the band, thus recovering the pattern for $U=0$, as reported in previous studies [46]. The objective of this study is to characterize this dielectric breakdown in terms of the electric field using the Landau-Zener formalism.
@@ 4 Threshold field in mesoscopic Mott insulators
Figure 4: Spectrum of the Landau-Zener Hamiltonian (Eq.(11)) for $v=1$. The plot on the right shows the gap at the avoided crossing, $\Delta E$, and $\epsilon_{\pm}(t)=\pm vt$ in dashed lines, ie. the energy levels for $\Delta E=0$.
@@
In these units, the threshold field is independent of $L$. However, this expression is only valid for finite lattices. In the thermodynamic limit, the spectrum of the Fermi-Hubbard model becomes flat and the LZ formalism is no longer applicable as $v\to 0$.
@@ 5.1 Non-equilibrium current oscillations
In order to excite the system to only a single state it is necessary to evolve the system no further than the first avoided crossing. Afterwards, the fraction of the wave function that has leaked to the excited state can leak again to the upper zones of the spectrum by transitions that are not within the scope of our theoretical formula. Then, to numerically study the dielectric breakdown we apply a linearly growing magnetic flux with a certain velocity from the ground state of $\Phi=0$ ($t=0$) until we reach the avoided crossing at $\Phi=\pi$ ($t=1/(2E)$). Once there, we maintain the magnetic flux constant (suppressing the electric field) and let the system evolve. Under the two-level approximation, the instantaneous wave function for $t>1/(2E)$ is
@@
Figure 5: Current response after switching of the field at the avoided crossing ($\Phi=\pi,\ t=1/2E$) for different on-site interactions. Different curves correspond to different fields. a) $U=1.5$. Solid line is $E=0.0078$, dashed $E=0.013$, and dotted $E=0.032$. b) $U=2.5$. Solid line is $E=0.045$, dashed $E=0.075$, and dotted $E=0.19$.
Furthermore, since the adiabatic ground (excited) state is at a maximum (minimum) of energy, the non-oscillatory terms in Eq. (20) must vanish according to Eq. (6), which causes the current to oscillate around zero. In Figure 5, the current of the system after switching off the field $(t>1/2E)$ shows an oscillatory pattern with a variable period and amplitude. As expected, the period depends only on the on-site interaction, whereas the amplitude of these oscillations grows with the field applied, evidencing a clear relation between the field and the movement of the electrons. Also, the oscillations are centered around zero as predicted. These findings strongly indicate that the two-level approximation is valid, and we also confirmed that the contribution of higher excited states is negligible using exact diagonalization.
These well-behaved oscillations allow us to obtain crucial information about equilibrium and non-equilibrium properties of the ring by studying their frequency and amplitude. The frequency equals the charge gap $\Delta$ of the spectrum, and to extract its value we compute the Fourier transform of the current after the field is switched off. For each interaction, which implies a particular energy gap and a single current frequency, we simulate a range of applied electric fields and average the set of obtained frequencies to gain statistical significance. The amplitude, which is field-dependent, contains information on the threshold field of the system, see next subsection. To obtain its magnitude, we take the maximum and minimum values of the current for each oscillation and do the average for all the oscillations obtained under the same conditions.
@@ 5.2 Finite coupling Landau-Zener
Note that although the evolution of the ground states does not start at an infinite time before the crossing, we expect this approximation to be valid for the cases in which the gap at the beginning of the evolution ($\Delta_{0}$) is much larger than the gap at the avoided crossing, $\Delta/\Delta_{0}\ll 1$. This condition does not hold when $\Delta\gg 1$ because the spectrum flattens, faster for greater L (see Eq. (10)). This results in $\Delta\approx\Delta_{0}$ which makes Eq. (24) no longer applicable.
@@ 6 Results
Figure 8: Threshold field $E_{th}(\Delta)$ for different number of sites. Dashed (dotted) line is Eq. (16) for $L=6$ ($L=10$). Empty markers show the $\xi<L$ regime. Data for $L=8$ is not shown for clarity but it is analogous to the other lattice sizes.
Figure 9: Threshold field $F_{th}(\Delta)$ for different ring sizes. Markers show the threshold field in the $\xi>L$ regime. Dashed line is Eq. (17), dotted line is the threshold field at thermodynamic limit $(F_{th}^{TL})$.
We have observed that the deviation of the data points in a particular gap $\Delta$ (and the corresponding $U$) in both lattice sizes coincides with the correlation length (which is a decreasing function of $U$) dropping below the lattice size. When $\xi<L$, the Drude weight starts to decay exponentially, implying a flattening of the spectrum and the breakdown of our approximation. In these conditions, the numerical fitting is no longer accurate, and the data deviates from the theoretical curve. To make this explicit, data points in this regime, $\xi<L$, are plotted with empty markers.
In Figure 9 we plot the threshold field in units of $F$ for $L=6,8,10$. We consider only the data points in the $\xi>L$ regime. Our theoretical formula in Eq. (17) is plotted in dashed lines, and, for comparison, the threshold field in the thermodynamic limit is plotted in dotted lines. The data points fall on our proposed theoretical curve, confirming that the threshold field is smaller for mesoscopic systems than in the thermodynamic limit. Furthermore, we plot three different system sizes to highlight the independence of the threshold field $(F_{th})$ on the system size.
@@ 7 Discussion
We study the dielectric breakdown of the mesoscopic 1D Fermi-Hubbard model. We propose a formula for the threshold field within the Landau-Zener formalism, which depends quadratically on the charge gap and is smaller than the threshold field at the bulk limit, in accordance to the finite-size effects of the charge stiffness. To check the validity of our prediction we numerically estimate the threshold field by studying the frequency and the amplitude of non-equilibrium current oscillations.
@@ Acknowledgements
This work has been funded by Grant PID2023-147475NB-I00 funded by MICIU/AEI/10.13039/501100011033 and FEDER, UE, by grants 2021SGR01095 and 2021SGR00907 from Generalitat de Catalunya, and by Project CEX2019-000918-M of ICCUB (Unidad de Excelencia María de Maeztu).

### v2

@@ 2.2 Linearly increasing magnetic flux
The non-equilibrium current is shown in Figure 3 with solid lines, while the persistent current is plotted in dotted lines. For $U=0$ the time-evolved wave function is at all times fully populating the same instantaneous eigenstate. As the eigenstate does not correspond, in general, to the ground state, the current does not show the Aharonov-Bohm period and shows no longer a sawtooth shape. In this case, it oscillates with the Bloch period and has higher amplitude than in equilibrium, as shown in Figure 3a).
Figure 3: Left column: Charge current of $L=10$ half-filled rings in the a) non-interacting $(U=0)$ and c) interacting $(U=2)$ regimes. Different curves correspond to different fields. The persistent current is displayed with dotted lines. A similar plot of the interacting regime can be found in [46]. Right column: b), d) Fourier transform of a) and c), respectively. In d), current curves without a clear frequency have been omitted for clarity.
For $U>0$, the situation is different. The existence of the charge gap allows the wave function to remain in the insulating state, as long as the evolution is adiabatic. That is why for sufficiently small fields the current is the same as in equilibrium, as shown in Figure 3c). For stronger electric fields, tunnelling to the excited spectrum becomes relevant, and the persistent current pattern is lost around the first avoided crossing, at $E\cdot t=0.5$. This is a signature of the dielectric breakdown of the system. It is relevant to note that the system does not necessarily excite to the first excited state, since there are excitations that do not couple to the electric field. For even larger fields, a Bloch oscillation pattern is observed, since the wave function can cross multiple avoided crossings and reach the top of the band, thus recovering the pattern for $U=0$, as reported in previous studies [46]. In Figure 3d) one can see how the frequency of the oscillations changes as the field increases: from the persistent current pattern (small period, high frequency) to the Bloch oscillation regime (large period, small frequency). The objective of this study is to characterize this dielectric breakdown in terms of the electric field using the Landau-Zener formalism.
@@ 4 Threshold field in mesoscopic Mott insulators
Figure 4: Spectrum of the Landau-Zener Hamiltonian (Eq. (11)) for $v=1$. The plot on the right shows the gap at the avoided crossing, $\Delta E$, and $\epsilon_{\pm}(t)=\pm vt$ in dashed lines, i.e., the energy levels for $\Delta E=0$.
@@
In these units, the threshold field is independent of $L$. We expect this equation to be reliable as long as the two-level approximation holds. However, when the work of the applied field becomes comparable to the gap of the next excited state, i.e., $2\pi E_{th}=LF_{th}\approx 2\Delta$, another state contributes to the wave function and the formula is no longer valid. In the lattice sizes considered in this work, this happens for $U\approx 5$. For lattice sizes up to $L\approx 10^{2}$, the two-level approximation remains valid for interaction strengths $0<U\approx 3$. In the thermodynamic limit, the spectrum of the Fermi-Hubbard model becomes flat and the LZ formalism is no longer applicable as $v\to 0$, so the formula is invalid regardless of the magnitude of the field.
@@ 5.1 Non-equilibrium current oscillations
In order to excite the system to only a single state it is necessary to evolve the system no further than the first avoided crossing. Afterwards, the fraction of the wave function that has leaked to the excited state can leak again to the upper zones of the spectrum by transitions that are not within the scope of our theoretical formula. Then, to numerically study the dielectric breakdown we apply a linearly growing magnetic flux with a certain velocity from the ground state of $\Phi=0$ ($t=0$) until we reach the avoided crossing at $\Phi=\pi$ ($t=1/(2E)$). Once there, we maintain the magnetic flux constant (suppressing the electric field) and let the system evolve. Under the two-level approximation, the instantaneous wave function for $t=1/(2E)$ is
@@
Figure 5: Left column: current response after switching off the field at the avoided crossing ($\Phi=\pi,\ t=1/2E$) for different on-site interactions. Different curves correspond to different fields. a) $U=1.5$. Solid line is $E=0.0078$, dashed $E=0.013$, and dotted $E=0.032$. c) $U=2.5$. Solid line is $E=0.045$, dashed $E=0.075$, and dotted $E=0.19$. Right column: b), d) Fourier transform of a) and c), respectively. The amplitudes are normalized by the peak. Line style of each field is conserved.
Furthermore, since the adiabatic ground (excited) state is at a maximum (minimum) of energy, the non-oscillatory terms in Eq. (20) must vanish according to Eq. (6), which causes the current to oscillate around zero. In the left column of Figure 5, the current of the system after switching off the field $(t>1/2E)$ shows an oscillatory pattern with a variable period and amplitude. As expected, the period depends only on the on-site interaction, whereas the amplitude of these oscillations grows with the field applied, evidencing a clear relation between the field and the movement of the electrons. Also, the oscillations are centered around zero as predicted. These findings strongly indicate that the two-level approximation is valid, and we also confirmed that the contribution of higher excited states is negligible using exact diagonalization.
These well-behaved oscillations allow us to obtain crucial information about equilibrium and non-equilibrium properties of the ring by studying their frequency and amplitude. The frequency equals the charge gap $\Delta$ of the spectrum, and to extract its value we compute the Fourier transform of the current after the field is switched off, see the right column of Figure 5. For each interaction, which implies a particular energy gap and a single current frequency, we simulate a range of applied electric fields and average the set of obtained frequencies to gain statistical significance. The amplitude, which is field-dependent, contains information on the threshold field of the system, see next subsection. To obtain its magnitude, we take the maximum and minimum values of the current for each oscillation and do the average for all the oscillations obtained under the same conditions.
@@ 5.2 Finite coupling Landau-Zener
We expect this approximation to be valid when 1) the two-level approximation holds, i.e. $2\pi E_{th}=LF_{th}<2\Delta$ which is true for fields up to $E\approx 1$ with lattices sizes of this work, and 2) when the gap at the beginning of the evolution ($\Delta_{0}$) is much larger than the gap at the avoided crossing, $\Delta/\Delta_{0}\ll 1$. This condition does not hold when $\Delta\gg 1$ because the spectrum flattens, faster for greater L (see Eq. (10)). This results in $\Delta\approx\Delta_{0}$ which makes Eq. (24) no longer applicable, regardless of the electric field applied.
@@ 6 Results
Figure 8: Threshold field $E_{th}(\Delta)$ for different number of sites. Dashed (dotted) line is Eq. (16) for $L=6$ ($L=10$). Empty markers show the $\xi<L{\color[rgb]{0,0,0}-1}$ regime. $\xi$ is estimated using Eq. (9). Data for $L=8$ is not shown for clarity but it is analogous to the other lattice sizes.
Figure 9: Threshold field $F_{th}(\Delta)$ for different ring sizes. Markers show the threshold field in the $\xi>L{\color[rgb]{0,0,0}-1}$ regime. Dashed line is Eq. (17), dotted line is the threshold field at the thermodynamic limit $(F_{th}^{TL})$.
We have observed that the deviation of the data points in a particular gap $\Delta$ (and the corresponding $U$) in both lattice sizes coincides with the correlation length (which is a decreasing function of $U$) dropping below the lattice size. When $\xi<L$, the Drude weight starts to decay exponentially, implying a flattening of the spectrum and the breakdown of our approximation. In these conditions, the numerical fitting is no longer accurate, and the data deviates from the theoretical curve. To make this explicit, data points in this regime, $\xi<L{\color[rgb]{0,0,0}-1}$, are plotted with empty markers.
In Figure 9 we plot the threshold field in units of $F$ for $L=6,8,10$. We consider only the data points in the $\xi>L{\color[rgb]{0,0,0}-1}$ regime. Our theoretical formula in Eq. (17) is plotted in dashed lines, and, for comparison, the threshold field in the thermodynamic limit is plotted in dotted lines. The data points fall on our proposed theoretical curve, confirming that the threshold field is smaller for mesoscopic systems than in the thermodynamic limit. Furthermore, we plot three different system sizes to highlight the independence of the threshold field $(F_{th})$ on the system size.
@@ 7 Discussion
We study the dielectric breakdown of the mesoscopic 1D Fermi-Hubbard model. We propose a formula for the threshold field within the Landau-Zener formalism, which depends quadratically on the charge gap and depends on the slope of the spectrum at the avoided crossing, analogously to single particle systems. Additionally, our formula is smaller than the threshold field at the bulk limit, in accordance to the finite-size effects of the charge stiffness. To check the validity of our prediction we numerically estimate the threshold field by studying the frequency and the amplitude of non-equilibrium current oscillations.
@@ Acknowledgements
This work has been funded by Grant PID2023-147475NB-I00 funded by MICIU/AEI/10.13039/501100011033 and FEDER, UE, by grants 2021SGR01095 and 2021SGR00907 from Generalitat de Catalunya, and by Project CEX2019-000918-M of ICCUB (Unidad de Excelencia María de Maeztu). T-G. J. acknowledges funding from the Spanish Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain, EU through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026.